A particle of mass 2 kg starts from rest on a inclined smooth surface with inclination 30 degree and length 2 metre after travelling along incline it is made to move on frictional horizontal surface the maximum distance it will travel before coming to rest on a frictional surface with frictional coefficient of 0.25 is

Grade:11

1 Answers

Vikas TU
7 years ago
After travelling 2 mter distance,
 
final velocity becomes as,
v^2 = 0^2 + 2gsin30*2
v = root(20) = m/s.
 
From friction sufrace starts, we would appyly WET,
Hence,
-(0.25*2*g*cos30)*x  = 0 –  0.5*2*v^2  { friction force is negative}
put v value and solve we get,
x = 2.312 meter.